Circular No. 7635

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE B. E. Schaefer, University of Texas at Austin, on behalf of the QUEST

collaboration (cf. IAUC 7387), reports on the discovery of eleven apparent supernovae discovered with the QUEST 16-CCD array camera on the Centro de Investigaciones de Astronomía 1-m Schmidt telescope at Llano del Hato, Venezuela. The total area searched was 264 deg^{2} between Mar. 25 and Apr. 1 to R = 20.8, with each clear night receiving independent drift scans through B, V, and R filters. Nightly scans continued until Apr. 4, so that the eleven objects all have well-sampled 3-color light curves over roughly three weeks. Each supernova was found by subtraction of QUEST reference images, and each new object has been confirmed via six or more independent images taken on three or more nights with subtractions from three independent reference images.

SN  2001 UT  α_{2000}  δ_{2000}  R  Offset
2001bu Mar. 25 10^{h}45^{m}22.^{s}19 − 1^{o}28′44′′.5 19.4 0′′ E, 0′′ N
2001bv Mar. 25 13 58 09.11 − 1 51 40.9 19.9 0′′ E, 2′′ S
2001bw Mar. 25 14 01 05.50 − 1 08 26.0 19.8 4′′ E, 4′′ N
2001bx Mar. 25 14 54 05.62 − 1 42 04.1 20.0 1′′ W, 3′′ N
2001by Mar. 25 15 25 52.96 − 2 10 28.7 20.3 0′′ E, 0′′ N
2001bz Mar. 25 15 30 13.76 − 0 09 26.0 20.7 2′′ W, 0′′ N
2001ca Mar. 25 15 35 41.15 − 0 16 21.6 19.4 0′′ E, 0′′ N
2001cb Mar. 25 15 38 44.08 − 0 22 56.6 20.5 0′′ E, 0′′ N
2001cc Mar. 25 15 44 13.61 − 1 53 15.5 20.0 0′′ E, 1′′ S
2001cd Mar. 25 15 46 38.83 − 0 23 09.2 19.9 1′′ W, 1′′ N
2001ce Mar. 25 16 39 28.49 − 1 27 17.9 17.9 0′′ E, 0′′ N

COMET P/2001 J1 (NEAT) As hinted on IAUC 7625, this is a short-period comet, and observa-

tions by C. W. Hergenrother, T. B. Spahr, and M. Nelson with the 1.8-m f/1 VATT Lennon telescope on May 27 make it clear that the orbital period is P ∼ 7.5–7.9 years. Spahr has also identified the comet with a very faint object (not described as cometary) discovered by A. E. Gleason with the Spacewatch telescope on 2000 Oct. 7 and placed on The NEO Confirmation Page but removed on Oct. 20 for lack of follow-up. The additional astrometry and orbital elements (P = 7.64 yr) are given on MPEC 2001-K43. S. Nakano has noted some rough similarity to the orbit of comet 3D/Biela.

2001 May 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green